OpenStudy (anonymous):

the simplified form of the radical expression. 4 times the square root of 8, all over the square root of 2. I'm not sure if it is 16 or 2 times the square root of 16.

OpenStudy (anonymous):

 Use: \[\frac{\sqrt{a}}{\sqrt{b}} = \sqrt{\frac{a}{b}}\]

OpenStudy (anonymous):

a = 8 and b = 2 there.. Put in the above formula and let us see what you get??

OpenStudy (campbell_st):

\[\frac{4\times2\times \sqrt{2}}{\sqrt{2}}\] should be easy to simplify

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so its 8?